OpenStudy (kaydra14):

Solve for w. P=2l + 2w

OpenStudy (anonymous):

w= width ^_^

OpenStudy (anonymous):

p=21+2w or, p-21 = 2w or, (p-21 )/2 = w that is the answer . r u getting this?

OpenStudy (anonymous):

@kaydra14

OpenStudy (kaydra14):

yes i got

OpenStudy (kaydra14):

can you help me wit another one please @Learner11

OpenStudy (anonymous):

sure ! But plz give me a medal

OpenStudy (kaydra14):

ok i will

OpenStudy (anonymous):

:)

OpenStudy (kaydra14):

Solve for the variable c in this equation a(b - c)=d

OpenStudy (kaydra14):

@Learner11

OpenStudy (anonymous):

here , a(b - c)=d or , ab -ac =d or , -ac= d -ab or , ac= ab-d or, c= (ab-d)/a got it ?